OpenStudy (anonymous):

One number is chosen from the list 2, 4, 6, and 8. Another is selected from the list 6, 7 and 8. Find the probability that their sum is odd.

OpenStudy (raden):

the events of odd sum are : (2,7),(4,7),(6,7),(8,7). number events = 4 the number of outcome = 4 * 3 = 12 so, P(odd sum) = 4/12 = 1/3